# JEL: C81, O32

### v1

@@
Across fields, the quantity of research has increased substantially, without an attendant increase in output. We argue that, in medicine, this indicator of declining productivity reflects a compositional shift toward low-capital, low-productivity research. Using a fine-tuned, open-source large language model, we construct a novel census of capital-intensive, high-productivity medical investment—clinical trials. Since 2010, the annual quantity of clinical trials has been constant. By contrast, the quantity of other forms of clinical research has increased substantially. Within clinical trials, there is substantial heterogeneity in productivity. The distribution of this heterogeneity is stable over time.
Keywords: Medical Innovation, Productivity, Large Language Models
@@ 1. Introduction
Technological progress in medicine is widely credited with enabling large improvements in human health (Easterlin,, 1999; Mokyr et al.,, 2002; Acemoglu and Johnson,, 2007). Breakthroughs at the turn of the century—the completion of the Human Genome Project, the advent of DNA sequencing, the adoption of combinatorial chemistry—promised transformation of diagnosis and treatment or, at least, an expansion of the set of available scientific opportunities. [^1] See Scannell et al., (2012) for details on these factors, in a broader discussion of declining productivity in the pharmaceutical industry. Combinatorial chemistry, in the 1980s and 1990s, increased the number of molecules (potential drug candidates) that a chemist could synthesize “perhaps 800-fold.” Similarly, since the first genome sequence was determined in the 1970s, DNA sequencing has become “over a billion times faster.” Yet, an emerging literature finds that biomedical research productivity is in a state of decline; see Goldin et al., (2024) for a review. This literature is structured around an incongruity identified by Bloom et al., (2020). The quantity of medical research—proxied by estimates of the number of clinical trials and scientific publications—has increased substantially, without an accompanying improvement in mortality. Many explanations center on characteristics of the “idea production function.” Perhaps low-hanging fruit have been depleted (the fishing out hypothesis, described in Jones,, 1995), [^2] See, also, Cowen, (2011) and Gordon, (2017). or else the educational costs to a young innovator of reaching the frontier of knowledge are increasingly high (the burden of knowledge hypothesis, due to Jones,, 2009). Though the specific parameter implicated differs, the mechanisms underlying these hypotheses suggest a structural change in research productivity.
This paper tests an alternative, compositional hypothesis. We document that increases in the quantity of medical research have been driven by increases in the quantity of unproductive, low-capital research with limited capacity to effect quantitatively significant changes to mortality. As a consequence, increased quantities of medical research need not indicate that the productivity of high-capital research investment is declining.
We construct a novel census of capital-intensive, high-productivity medical investment—clinical trials. We contrast this set with its complement in clinical research, proxied by publications that cite clinical trials. This comparison is natural and reflects the standard hierarchy of medical evidence (Jones and Podolsky,, 2015). Decision-making in medicine—by regulators, firms, insurers, physicians, and patients—hinges on the causal, mortality-relevant data produced by costly, time-intensive randomized experiments (Chavez-MacGregor and Giordano,, 2016; Alsan et al.,, 2024). Although research engaging with trial evidence undoubtedly produces valuable information, it does so at comparatively low cost and with lesser impact (Kones et al.,, 2014). Since 2010, we find that the quantity, quality, and composition of published clinical trials has been stable. By contrast, the quantity of other clinical research has doubled.
No existing dataset contains sufficient information to assess the composition of clinical research. Administrative databases—such as ClinicalTrials.gov—aim to provide an index of trials, but suffer from widespread noncompliance with reporting requirements (DeVito et al.,, 2020). Proprietary databases take as an input the contents of ClinicalTrials.gov and other registries and, thus, capture changes in reporting patterns over time (e.g., Cortellis,, 2024). Best practices for identifying trials in publication data do so imprecisely, with both high true and false positive rates (Thomas et al.,, 2021).
@@
We construct a new census of clinical trials. We collect all records of scientific publications indexed in the National Library of Medicine’s PubMed / MEDLINE database, from 2010 forward, that disclose the results of a clinical trial studying a medicine in human subjects. Our definition is restrictive and intended to capture only those research inputs that are most relevant for human mortality. Our goal is to identify clinical trials in our sample of interest—from the set of roughly 35 million PubMed records—with an accuracy and precision close to that of a human labeller. Standard machine learning methods perform poorly and hand-labelling at scale is infeasible. Instead, we train a bespoke large language model, optimized for our task.
@@
We proceed in stages. We devise an interface to iteratively revise model prompts. With the highest-performing prompts, we use proprietary language models, OpenAI’s GPT-3.5 and GPT-4, to classify a large number of records (Achiam et al.,, 2023; Bubeck et al.,, 2023). These noisy labels are used as training data to fine-tune a set of open-source models. The resulting model matches the performance of the best proprietary model and produces a sample of approximately 150,000 clinical trials.
@@
We document two sets of facts. First, there is quantitatively significant heterogeneity in both the composition of medical research and the composition of clinical trials. While clinical trial counts are essentially constant over time, the quantity of papers citing trials has increased by nearly a factor of two. This trend reflects a large increase in the quantity of research originating from China, in addition to other countries who, individually, do not account for a large share of the total quantity of medical research. Within clinical trials, proxies for quality and informativeness suggest considerable heterogeneity. Roughly half of trials in our sample are never cited by a leading medical journal and nearly 15 percent are never cited by another scientific publication. Second, these forms of heterogeneity within trials are stable over time. We find that—across measures of quality, importance, and geography—the distribution of clinical trial heterogeneity is unchanging. Taken together, we interpret these facts as evidence that the productivity of high-capital medical investment is stable over our time period.
Our choice of setting reflects two substantive considerations. First, the determinants of technological progress in medicine are interesting in their own right. [^3] Cutler and McClellan, (2001) observe that technological change has accounted for the bulk of increases in health spending over time, but that health has improved alongside these increases. In a series of case studies, they argue that the benefits of technological progress exceed their costs. See, also, Almond et al., (2010), who estimate that the returns to medical spending are positive on the margin. Second, and more centrally, nearly all micro-level evidence on rising research investment, and declining productivity, comes from studies of clinical trials and scientific publications. [^4] See Goldin et al., (2024) for a summary. Notable exceptions are studies of agriculture and semiconductors. Measured increases in the quantity of clinical trials (Bloom et al.,, 2020) and average “declining disruptiveness” of scientific papers (Park et al.,, 2023) are at the center of active debates about whether research in science and medicine has hit a point of diminishing returns (e.g., Dieppe and Kose,, 2020). [^5] See discussions in the popular press, e.g. Collison and Nielsen, (2018); Broad, (2023); Thompson, (2021); Piper, (2023). Concerns about declining productivity have influenced a wave of policy interest in new mechanisms to promote research and development, as well as the revival of “industrial policy” (see e.g., Council of Economic Advisers,, 2021, 2024; Lindsey and Hammond,, 2020; Hausmann,, 2023). A precise characterization of compositional changes in this industry, then, is a directly policy-relevant object.
We propose and validate the existence of quantitatively significant heterogeneity in medical research. [^6] Park et al., (2023) examine “disruptiveness” in scientific research, using data on the universe of publications across six decades. They document that the absolute counts of highly disruptive papers remain constant, consistent with our finding that high-capital, high-productivity research remains stable. Our setting allows for a clear delineation between types of research investments and, importantly, allows us to take a stand on the value of information produced by these distinct types for welfare-relevant outcomes. This is the primary contribution of this paper. We measure research inputs with a precision that is new to this literature, enabled by the existence of novel technology for data construction at scale (Achiam et al.,, 2023; Bubeck et al.,, 2023). Our decomposition suggests that measures of research investment that neglect these differences—and, instead, infer uniform productivity or capital for scientific investments—can yield misleading conclusions about industry-wide productivity.
@@
This paper contributes to a growing literature on the causes of declining productivity, in general and in medicine. Additional explanations include market imperfections and failures—including barriers to innovation (see e.g., Aghion et al.,, 2019), inadequate public sector support for research (Gruber and Johnson,, 2019), and productivity imbalances across sectors (Acemoglu et al.,, 2024). Our work is related to a literature—which Syverson, (2017) summarizes as putting forth a “mismeasurement hypothesis”—that views the measured decline in aggregate productivity as illusory (see e.g., Mokyr,, 2014). In contrast, our hypothesis is not that productivity is mismeasured, per se, but that failure to account for heterogeneity can generate a misattribution of trends. More directly, Myers and Pauly, (2019), Scannell et al., (2012), and Cockburn, (2006) investigate the determinants of declining productivity in the pharmaceutical sector, with a focus on large increases in factor prices for clinical trials rather than quantities.
Finally, we draw on, and contribute to, a long literature on the measurement of innovation. Since Pakes, (1986), economists have recognized that patents—the most commonly used measure of research output—differ greatly in economic and technological significance. Griliches, (1990) provides a classic discussion of the difficulties associated with using patents as economic indicators, given this heterogeneity. [^7] Bryan and Williams, (2021) provide an updated survey and discuss the challenges of devising credible measures of innovative effort. Our work builds on these insights to investigate the implications of heterogeneity in inputs for the use of medical research as a productivity indicator.
2. Productivity, Capital, and Medical Research
@@
Studies of research productivity have, at their core, some notion of an “idea production function”—a mapping from research inputs (effort) to outputs (ideas). The canonical Romer, (1990) specification is given by

### v2

@@
A recent literature contends that the returns to research in medicine are declining. This conclusion is premised on two empirical findings: the quantity of medical research is rising, while improvements in health outcomes are stagnant. We argue that the first finding is an artifact of conceptual and empirical problems in the measurement of research. We propose a method for annotating unstructured text based on fine-tuning a large language model. We apply this method to construct a census of published clinical trials. We argue that clinical trials function as a natural proxy for productivity-relevant research in medicine. In our data, the quantity, quality, and composition of clinical trials have been stable since 2010. We find that measured increases in the total quantity of medical research are driven by shifts in the composition of scientific publications that do not correspond to real increases in the quantity of productivity-relevant research. We document large increases in the growth of non-trial publications, driven by publications from China and publications that review existing literature. This increase in quantity is coincident with a decline in various measures of average publication quality. Taken together, we find no support for declining medical research productivity over this period.
Keywords: Medical Innovation, Productivity, Clinical Trials, Large Language Models, Generative AI
@@ 1. Introduction
Two stylized facts support the argument that the returns to medical research are diminishing. [^1] Goldin et al., (2024) survey evidence related to declining productivity across the economy, including evidence of diminishing returns to research investment in sectors beyond medicine. The quantity of resources directed toward medical research, measured in dollars and scientific publications, appears to be increasing (Scannell et al.,, 2012; DiMasi et al.,, 2016; Bloom et al.,, 2020; Park et al.,, 2023). [^2] See Park et al., (2023), Bloom et al., (2020), and Scannell et al., (2012) for case studies examining the quantity of investment in science, medicine, and drug development, respectively. Bloom et al., (2020) collect case studies from other research-intensive fields. See, also, Chu and Evans, (2021), who document large increases in the quantity of papers in certain scientific fields and examine associated patterns in citation networks. Meanwhile, improvements in human health (e.g., increases in life expectancy) continue at the same steady pace (Mehta et al.,, 2020; Ho and Hendi,, 2018). If more resources are necessary to yield the same rates of technological progress, simple models of “idea production” imply that research productivity—the efficiency with which research is converted into new and useful knowledge—must be falling (Jones,, 1995; Bloom et al.,, 2020). [^3] In some models, these patterns can be reconciled without diminishing returns to research. For instance, Young, (1998) investigates the idea that investment in new varieties of existing products can yield the same trends.
These facts, and their forecast of technological stagnation in a sector closely linked to health and welfare (Cutler and McClellan,, 2001; Mokyr et al.,, 2002; Acemoglu and Johnson,, 2007), have influenced prominent policy debates. [^4] These facts were most prominently documented in Bloom et al., (2020). Concerns about technological stagnation in modern policy discourse predate Bloom et al., (2020) and respond in part to Cowen, (2011) and Gordon, (2017). As we discuss below, the prospect of diminishing returns to scientific research has long been a subject of concern. Proposals that would restructure public subsidies administered by the National Institutes for Health (NIH) (e.g., Executive Office of the President,, 2019; Moss et al.,, 2020; The Economist,, 2022), relax the regulatory standards for approval of new medicines by the U.S. Food and Drug Administration (FDA) (e.g., Roy,, 2021; Chertman,, 2023; Graboyes and Reimschisel,, 2017), and prevent the Centers for Medicare and Medicaid Services (CMS) from negotiating drug prices (e.g., York,, 2023; Long,, 2023; Inflation Reduction Act,, 2022) are premised, in no small part, on the concern that science and medicine are in a state of decline. [^5] This concern has been explored extensively in the popular press. Collison and Nielsen, (2018), Broad, (2023), Thompson, (2021), and Piper, (2023), among others, interpret these facts in the popular press as a “warning sign” for the future of technological progress (Piper,, 2023).
In this paper, we find that one of these stylized facts—the increase in the quantity of resources directed toward medical research—is an artifact of conceptual and empirical problems in the measurement of medical research. An ideal, and infeasible, measure of the quantity of research captures the amount of “effort” directed toward increasing the stock of useful (medical) knowledge. Informative proxies for this measure have two characteristics. First, as researchers differ in skill and research designs differ in quality, the proxy must account for heterogeneity in the likelihood that research projects yield “productivity-relevant” information. Second, units of the proxy must be comparable over time. If the quantity of research embedded in, say, a patent or paper decreases with time, increases in the quantity of patents or papers provide very little insight into research productivity.
We find that existing measures of medical research fall short on both accounts. The primary contribution of this paper is the construction of a measure of medical research—a census of published clinical trials—that addresses both issues and fills this gap. We show that standard data construction tools classify research imprecisely, and that this imprecision yields inaccurate conclusions about trends in productivity. Thus, we introduce a method for annotating unstructured text at scale. When applied to a corpus of scientific text, this method allows us to identify and classify heterogeneous forms of research, relaxing a long-standing data constraint in the economics of innovation.
We fine-tune an open-source large language model and devise a workflow that allows us to identify clinical trial records from the unstructured text of scientific publications with an accuracy comparable to that of a human labeler. In our data, the quantity, quality, and composition of clinical trials have been stable since 2010. We find evidence that measured increases in the quantity of scientific publications capture shifts in scientific incentives that are not indicative of real increases in the quantity of (productivity-relevant) research. Taken together, we find no support for claims of declining medical research productivity over this period.
We begin by arguing that clinical trials are a natural proxy for the research object of interest in this setting. Trials are discrete, regulated experiments that generate useful statistical information about new and existing technologies (Chavez-MacGregor and Giordano,, 2016; Alsan et al.,, 2024). These experiments are designed in consultation with regulators after sufficiently robust preliminary evidence has been produced but before health technologies can be marketed. Trials top the hierarchy of medical evidence employed by firms, regulators, physicians, and patients (Jones and Podolsky,, 2015). Since the early 2000s, clinical trials have been reported in a consistent, standardized format across scientific publications. [^6] In the 1990s and 2000s, medical experts created a checklist that provides authors of papers reporting the results of clinical trials with a standardized structure and list of elements for publication. The most recent Consolidated Standards of Reporting Trials (CONSORT) Statement was published in 2010 and is now the reporting standard for most scientific journals that publish the results of trials. See Laine et al., (2007); Schulz et al., (2010) for details. To our knowledge, this type of standardization is unique to this form of science.
No existing dataset allows us to track quantities of clinical trials over time. Although dollar-denominated expenditures are an obvious metric, available records of investment in trials are self-reported, in some form, by pharmaceutical firms and are essentially impossible to validate independently. Administrative databases—such as ClinicalTrials.gov—aim to provide an index of trials, but suffer from widespread noncompliance with reporting requirements (e.g., DeVito et al.,, 2020). Proprietary databases (e.g., Cortellis,, 2024) take as an input the contents of publicly-maintained registries and, thus, are biased by these same noncompliance patterns.
@@
We construct a new census of clinical trials indexed in scientific publications. We collect all records in the National Library of Medicine’s PubMed / MEDLINE database, from 2010 forward, that disclose the results of a clinical trial that studied the effects of a medicine in human subjects. This definition is intentionally restrictive and aims to include only those records that clearly report trials. Our goal is to identify clinical trials in our sample of interest—from the set of roughly 35 million PubMed records—with an accuracy and precision close to that of a human labeler.
Manually identifying clinical trial records, at this scale, is infeasible. [^7] A small number of papers, including Oostrom, (2024), examine medical investment using small samples of clinical trial records that are constructed by hand. Standard methods perform poorly. In particular, we show that documented increases in the quantity of clinical trials during this time period, as reported by Bloom et al., (2020), are driven by a categorization error caused by keyword-based searches. [^8] Specifically, Bloom et al., (2020) categorize many records as clinical trials that discuss trials (e.g., meta-analyses and literature reviews). We find that changes in the scientific publishing landscape, discussed in Section 4, increase the size of this categorization error over time. In our period of interest, this generates a spurious increase in the measured quantity of trials. Machine learning methods used in the health services literature similarly lack the precision necessary for our purposes (Thomas et al.,, 2021). Instead, we leverage recent advances in generative AI, which allow for data extraction from unstructured text at scale. We devise a workflow for extracting precisely defined objects from a large corpus of text, and train a bespoke large language model that is specifically optimized for this task.
@@
We proceed in stages. We devise an interface to iteratively revise model prompts. With the highest-performing prompts, we use proprietary language models, OpenAI’s GPT-3.5 and GPT-4, to classify a large number of records (Achiam et al.,, 2023; Bubeck et al.,, 2023). These noisy labels are used as training data to fine-tune a set of open-source models. An ensemble of the fine-tuned models matches the performance of the best proprietary model, with false positive and false negative rates below 5 percent, and produces a sample of approximately 150,000 publications that report the results of clinical trials from 2010 to 2022.
@@
These data allow us to document two sets of facts about the quantity of medical research over time. First, we find that the quantity of clinical trials has been effectively constant, year-on-year, across our period of interest. We collect a set of ex-ante and ex-post quality measures, derived from scientific publication data. On these dimensions, we find considerable heterogeneity. Roughly half of clinical trial publications in our sample are never cited by a leading medical journal. Nearly 15 percent are never cited by any other scientific publication. However, we find that the distribution across clinical trials of measures of quality, importance, and geography is unchanging. Taken together, we interpret these patterns as evidence that “productivity-relevant” research in medicine is essentially constant in quantity and stable in composition over our time period. This contradicts existing findings, at least for our period of interest, and calls into question the conclusion that the productivity of medical research is falling.
Second, our data allow us to examine the source of measured increases in the quantity of scientific publications (Bloom et al.,, 2020; Park et al.,, 2023; Chu and Evans,, 2021). We contrast clinical trial production with trends in the publication of other forms of medical research. While the quantity of clinical trials is constant across our period of interest, the quantity of other medical research—defined, in our preferred version, as the set of papers citing clinical trials—has increased by nearly a factor of two. [^9] In Section 4, we inspect the contents of this set and document that nearly all records fall into one of three categories: observational studies, preclinical studies, and meta-analyses/literature reviews. In doing so, we confirm that our focus on clinical trials is appropriate. These increases in quantity have coincided with shifts in the quality and composition of these other types of research. We highlight three patterns. First, the large increase in the total number of medical papers in our sample is largely driven by researchers in China. Second, this rise in quantity coincides with a decrease (on various measures) in the average quality of publications. Third, there is a substantial increase in the quantity of meta-analyses and literature reviews during our period of interest. [^10] These patterns are consistent with the geometric increase in the quantity of meta-analyses produced by Chinese researchers, documented in Ioannidis et al., (2013). We interpret these patterns as evidence that—for this category of other medical research—the quantity of productivity-relevant research “embedded” in a single document may be declining over time. Quantities of publications, then, provide essentially no insight on the question of research productivity. [^11] Park et al., (2023) examine the “disruptiveness” of scientific research, using data on the universe of publications across six decades. Their findings are similar to ours, in the sense that they document that the absolute counts of highly disruptive papers remain constant over time, while average “disruptiveness” is declining. They interpret this as evidence that science is becoming less disruptive over time. Our paper differs in two important ways. First, we distinguish between types of publications on the basis of the economic value they produce. Second, we document that measured increases in the quantity of publications are disproportionately concentrated in types of publication that focus on diffusion, rather than production, of new information. Pooled averages, then, may be difficult to interpret.
@@
Our findings speak directly to ongoing policy discussions and branches of the macroeconomic literature, which consider the hypothesis that the rate of technological progress may be declining. Nearly all micro-level evidence on rising levels of research investment comes from studies of scientific publications and, in particular, clinical trials. [^12] See Goldin et al., (2024) for a summary. Notable exceptions are studies of agriculture and semiconductors. We show that these findings result from conceptual and empirical challenges in the quantification of innovation. Thus, our work is most closely related to a series of papers—which Syverson, (2017) summarizes as putting forth a “mismeasurement hypothesis”—that view the measured decline in aggregate productivity as illusory (see e.g., Mokyr,, 2014). [^13] Additional explanations include market imperfections and failures—including barriers to innovation (see e.g., Aghion et al.,, 2019), inadequate public sector support for research (Gruber and Johnson,, 2019), and productivity imbalances across sectors (Acemoglu et al.,, 2024). More directly, Myers and Pauly, (2019), Scannell et al., (2012), and Cockburn, (2006) investigate the determinants of declining productivity in the pharmaceutical sector, with a focus on large increases in factor prices for clinical trials rather than quantities.
In doing so, we draw on, and contribute to, a long literature on the measurement of invention. Pakes and Griliches, (1980) begin a discussion of inventive activity led by U.S. firms by remarking: “Perhaps the most serious task facing empirical work in the area of ‘technological change’ and ‘invention and innovation’ is the construction and interpretation of measures (indices) of advances in knowledge.” In a more recent survey, Bryan and Williams, (2021) note that, although much progress has been made over half a century of research, construction of such measures remains an open challenge. Decades of work have focused, in particular, on the development of proxies for the outputs of technological innovation. An especially prominent approach, due to Hall et al., (2000), uses citation-weighted counts of patents. Kogan et al., (2017) and Kelly et al., (2021) provide two recent, novel, alternatives. Our focus, by contrast, is on the difficulties associated with measuring the inputs of innovation, which can similarly bias estimates of macroeconomic trends. While scientific papers are a natural alternative to patents on the input side, existing methods for the use of patent data are not easily applied to publications (Popp,, 2015; Popp et al.,, 2010). [^14] In particular, citations across scientific papers are not analogous to citations across patents. Whereas patent citations are legal objects, which are defined in part by patent examiners and delineate the boundaries of a right enforceable by courts, citations across scientific papers reflect norms and networks. See Hall et al., (2000).
The idea that patterns like those that motivate this paper—rising research quantities and stagnant outcomes—signal the “exhaustion of inventive opportunities” (a phrase due to Griliches, (1994)) is, similarly, an old one. Early appearances in the literature include Robert K. Merton’s 1935 discussion of the “possibility of a slackening in the rate of technologic advance” (Merton,, 1935), Alfred B. Stafford’s 1952 paper titled “Is the Rate of Invention Declining?” (Stafford,, 1952) and Jacob Schmookler’s various investigations (Schmookler,, 1952, 1954, 1966) into patterns suggestive of diminishing returns to research. Our reconciliation of these patterns hews closely to that of this earlier literature, which viewed the evidence base on productivity decline as inconclusive, given the shortcomings of available measures of innovation (Pakes and Griliches,, 1980; Griliches,, 1990, 1994). [^15] Griliches, (1990) summarizes his version of this conclusion more prosaically: “One can always worry that the world is coming to an end. Someday it undoubtedly will, but it does not look as if the end is already upon us, at least not yet.” This paper differs on one important dimension. We show that advances in data construction technology, if deployed carefully, provide researchers with sufficiently precise measures to draw conclusions.
A nascent literature investigates the use of large language models in economics (Dell,, 2024; Korinek,, 2023). We join a small set of papers, e.g., Bartik et al., (2023) and Dell, (2024), that use generative AI as a tool to construct high-quality economic data. Our approach is distinct in this literature, in that we fine-tune an open-source language model using data extracted from a larger, proprietary model. In doing so, we show that it is feasible for researchers to achieve the quality of hand-constructed data at a practically interesting scale, with a reproducible and transparent methodology, at a feasible cost. Our use of these technologies is distinct from other applications considered in e.g., Korinek, (2023), Manning et al., (2024), Bybee, (2023), Horton, (2023), and Mullainathan and Rambachan, (2024), which focus on the use of generative AI for other aspects of the scientific process, including hypothesis identification, survey participation, idea generation, and scientific writing.
More generally, we contribute to the large empirical literature that uses text as data (Gentzkow et al.,, 2019). We anticipate that our procedure for the construction of new, precise data on scientific research may be applicable to other contexts where unstructured text is available and manual construction of data is infeasible. We highlight a set of steps—in validation, prompt design, and fine-tuning of open source models—that yield data of sufficiently high quality at comparatively low cost.
2. Measuring Research
@@
Studies of research productivity have, at their core, some notion of an “idea production function”—a mapping from research (e.g., scientists, dollars, research capital) to new, useful information. The canonical Romer, (1990) specification, which motivates a recent empirical investigation in Bloom et al., (2020), is given by